# Another algorithm template

### v11

@@ 3.2 Interpolation approximation
We now consider the case $\alpha>1/q_{2}-1/2$. Again, denote by $A(\xi)$ the sum taken over $(k,{\boldsymbol{s}})\not\in G(\xi)$ in the right-hand side of (3.24). We split $A(\xi)$ into two sums as
$A(\xi)\ =\ \sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi,\ \sigma_{2;{\boldsymbol{s}}}\leq 2^{-\tau k}\xi^{\vartheta}}\delta_{k}(v_{\boldsymbol{s}})\,I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}\ +\sum_{\sigma_{2;{\boldsymbol{s}}}>2^{-\tau k}\xi^{\vartheta}}\delta_{k}(v_{\boldsymbol{s}})\,I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}.$
Denote by $A_{1}(\xi)$ and $A_{2}(\xi)$ the sums in the right-side hand. From (3.24) we derive
$\big{\|}v-{\mathcal{I}}_{G(\xi)}v\big{\|}_{{\mathcal{L}}_{2}(X^{1})}\ \leq\ \big{\|}v-{\mathcal{S}}_{G(\xi)}v\big{\|}_{{\mathcal{L}}_{2}(X^{1})}+\|A_{1}(\xi)\|_{{\mathcal{L}}_{2}(X^{1})}+\|A_{2}(\xi)\|_{{\mathcal{L}}_{2}(X^{1})}.$ (3.32)
Lemma 3.5 gives
@@
$\begin{split}\|A_{1}(\xi)\|_{{\mathcal{L}}_{2}(X^{1})}\ &\leq\ \sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi,\ \sigma_{2;{\boldsymbol{s}}}\leq 2^{-\tau k}\xi^{\vartheta}}\|\delta_{k}(v_{\boldsymbol{s}})\|_{X^{1}}\big{\|}I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}\big{\|}_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\\ \ &\leq\ C\,\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi,\ \sigma_{2;{\boldsymbol{s}}}\leq 2^{-\tau k}\xi^{\vartheta}}\|v_{\boldsymbol{s}}\|_{X^{1}}\big{\|}I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}\big{\|}_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\\ \ &\leq\ C\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}\|v_{\boldsymbol{s}}\|_{X^{1}}\sum_{k=0}^{k(\xi)}\big{\|}I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}\big{\|}_{L_{2}({\mathbb{R}}^{\infty},\gamma)}.\end{split}$
Similarly to (3.30), we have
$\big{\|}I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}\big{\|}_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\ \leq\ p_{\boldsymbol{s}}(\theta,\lambda)$
with the same $\theta$ and $\lambda$ as in (3.19). This gives the estimate
$\begin{split}\|A_{1}(\xi)\|_{{\mathcal{L}}_{2}(X^{1})}\ &\leq\ C\log\xi\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}\|v_{\boldsymbol{s}}\|_{X^{1}}\,p_{\boldsymbol{s}}(\theta,\lambda)\,\\ \ &\leq\ C\,\log\xi\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}\|v_{\boldsymbol{s}}\|_{X^{1}}\,p_{\boldsymbol{s}}(\theta,\lambda).\end{split}$
@@
$\begin{split}\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}\|v_{\boldsymbol{s}}\|_{X^{1}}\,p_{\boldsymbol{s}}(\theta,\lambda)\ &\leq\ \left(\sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}(\sigma_{1;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{1}})^{2}\right)^{1/2}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}p_{\boldsymbol{s}}^{2}(\theta,\lambda)\sigma_{1;{\boldsymbol{s}}}^{-2}\right)^{1/2}\\[6.45831pt] \ &\leq\ C\,\left(\sum_{\sigma_{1;{\boldsymbol{s}}}>\xi^{1/q_{1}}}p_{\boldsymbol{s}}^{2}(\theta,\lambda)\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\sigma_{1;{\boldsymbol{s}}}^{-(2-q_{1})}\right)^{1/2}\\[6.45831pt] \ &\leq\ C\xi^{-(1/q_{1}-1/2)}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}p_{\boldsymbol{s}}(2\theta,\lambda)\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\right)^{1/2}\leq C\xi^{-(1/q_{1}-1/2)}.\end{split}$
This means that
$\|A_{1}(\xi)\|_{{\mathcal{L}}_{2}(X^{1})}\ \leq\ C\xi^{-(1/q_{1}-1/2)}\log\xi.$ (3.34)
Observe that $\vartheta\alpha/\tau=1/q_{1}-1/2$ and $\alpha/\tau=1-q_{2}/2$. Hence, employing (2.4), the assumption (3.18) and the Hölder inequality we get the estimates for $\|A_{2}(\xi)\|_{{\mathcal{L}}_{2}(X^{1})}$:
$\begin{split}\|A_{2}(\xi)\|_{{\mathcal{L}}_{2}(X^{1})}\ &\leq\ \sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\ \sum_{2^{k}>\left(\xi^{\vartheta}\sigma_{2;{\boldsymbol{s}}}^{-1}\right)^{1/\tau}}\|\delta_{k}(v_{\boldsymbol{s}})\|_{X^{1}}\,\big{\|}I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}\big{\|}_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\\[6.45831pt] \ &\leq\ C\,\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\ \sum_{2^{k}>\left(\xi^{\vartheta}\sigma_{2;{\boldsymbol{s}}}^{-1}\right)^{1/\tau}}2^{-\alpha k}\|v_{\boldsymbol{s}}\|_{X^{2}}\,p_{\boldsymbol{s}}(\theta,\lambda)\\[6.45831pt] \ &\leq\ C\,\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\ \big{(}\sigma_{2;{\boldsymbol{s}}}^{-1/\tau}\xi^{\vartheta/\tau}\big{)}^{-\alpha}\|v_{\boldsymbol{s}}\|_{X^{2}}\,p_{\boldsymbol{s}}(\theta,\lambda)\\[6.45831pt] \ &\leq\ C\,\xi^{-\vartheta\alpha/\tau}\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\ \sigma_{2;{\boldsymbol{s}}}^{\alpha/\tau}\|v_{\boldsymbol{s}}\|_{X^{2}}\,p_{\boldsymbol{s}}(\theta,\lambda)\\[6.45831pt] \ &\leq\ C\,\xi^{-(1/q_{1}-1/2)}\ \sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\ \sigma_{2;{\boldsymbol{s}}}^{1-q_{2}/2}\|v_{\boldsymbol{s}}\|_{X^{2}}\,p_{\boldsymbol{s}}(\theta,\lambda)\\ \ &\leq\ C\,\xi^{-(1/q_{1}-1/2)}\ \left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}(\sigma_{2;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{2}})^{2}\right)^{1/2}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}p_{\boldsymbol{s}}^{2}(\theta,\lambda)\sigma_{2;{\boldsymbol{s}}}^{-2}\right)^{1/2}\\ \ &\leq\ C\,\xi^{-(1/q_{1}-1/2)}.\end{split}$
@@
$\|A_{2}(\xi)\|_{{\mathcal{L}}_{2}(X^{1})}\ \leq\ C\xi^{-(1/q_{1}-1/2)}.$ (3.35)
Combining (3.32)–(3.35) leads to the estimate
@@
We give estimate the dimension of the space ${\mathcal{V}}(G(\xi))$. Put $q:=\tau q_{2}$ and define $q^{\prime}$ by $1/q^{\prime}+1/q=1$. Since $\alpha>1/q_{2}-1/2$, we have that $q>1$, consequently, by using the Hölder inequality and (3.18), we derive that
@@
$\xi_{n}^{-(1/q_{1}-1/2)}\ \leq\ (2M)^{\frac{\alpha}{\alpha+\delta}}\,n^{-(1/q_{1}-1/2)\frac{\alpha}{\alpha+\delta}}.$
@@
where $k_{n}:=\lfloor\log_{2}\xi_{n}\rfloor$ and for $k\in{\mathbb{N}}_{0}$ and $\xi>1$,
@@
Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that for the operator ${\mathcal{I}}^{*}_{G_{\rm ev}(\xi_{n})}:{\mathcal{L}}_{2}^{\mathcal{E}}(X^{2})\to{\mathcal{V}}(G_{\rm ev}(\xi_{n}))$, we have that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
@@
Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda(\xi_{n}))|\leq n$ and
@@
Then there exists a constant $C$ such that for each $m\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
@@ 4 Integration
    Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
@@
    Let $\phi\in(X^{1})^{\prime}$ be a bounded linear functional on $X^{1}$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (4.5)
@@
$\begin{split}\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v_{\rm ev}({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\langle\phi,v\rangle\right|\ &\leq\ \int_{{\mathbb{R}}^{\infty}}\left|\langle\phi,v_{\rm ev}({\boldsymbol{y}})-{\mathcal{I}}^{*}_{G_{\rm ev}(\xi_{n})}v_{\rm ev}({\boldsymbol{y}})\rangle\right|\,{\rm d}\gamma({\boldsymbol{y}})\\[6.45831pt] \ &\leq\ \int_{{\mathbb{R}}^{\infty}}\|\phi\|_{(X^{1})^{\prime}}\|v_{\rm ev}({\boldsymbol{y}})-{\mathcal{I}}^{*}_{G_{\rm ev}(\xi_{n})}v_{\rm ev}({\boldsymbol{y}})\|_{X^{1}}\,{\rm d}\gamma({\boldsymbol{y}})\\[6.45831pt] \ &\leq\ C\|v_{\rm ev}-{\mathcal{I}}^{*}_{G_{\rm ev}(\xi_{n})}v_{\rm ev}\|_{{\mathcal{L}}_{1}(X^{1})}\\[6.45831pt] \ &\leq\ C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}\end{split}$
@@
    Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
@@
    Let $\phi\in X^{\prime}$ be a bounded linear functional on $X$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$ there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}(\xi_{n})}\langle\phi,v\rangle\right|\ \leq\ Cn^{-(1/q-1/2)}.$ (4.8)
The constants $C$ in (4.7) and (4.8) are independent of $v$ and $n$.
@@
Under the hypothesis of Theorem 3.1, assume additionally that the sequences $Y_{m}$, $m\in{\mathbb{N}}_{0}$, are symmetric. Define for $\xi>1$
$G_{2}(\xi):=\ \begin{cases}\big{\{}(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{2}:\,2^{k}\sigma_{2;{\boldsymbol{s}}}^{q_{2}}\leq\xi\big{\}}\quad&{\rm if}\ \alpha\leq 1/q_{2}-1/2;\\ \big{\{}(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{2}:\,\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi,\ 2^{\tau k}\sigma_{2;{\boldsymbol{s}}}\leq\xi^{\vartheta}\big{\}}\quad&{\rm if}\ \alpha>1/q_{2}-1/2,\end{cases}$ (4.9)
@@
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{{\mathcal{Q}}}_{G_{2}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (4.11)
@@
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{Q}_{\Lambda_{2}(\xi_{n})}\langle\phi,v\rangle\right|\ \leq\ Cn^{-(1/q-1/2)}.$ (4.14)

### v14

@@ 3.2 Interpolation approximation
We now consider the case $\alpha>1/q_{2}-1/2$. Lemma 3.5 gives
@@
We split $A(\xi)$ into two sums as
$A(\xi)=A_{1}(\xi)+A_{2}(\xi),$
where
$A_{1}(\xi):=\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi,\ \sigma_{2;{\boldsymbol{s}}}\leq 2^{-\tau k}\xi^{\vartheta}}\delta_{k}(v_{\boldsymbol{s}})\,p_{\boldsymbol{s}}(\theta,\lambda),$
and
$A_{2}(\xi):=\sum_{\sigma_{2;{\boldsymbol{s}}}>2^{-\tau k}\xi^{\vartheta}}\delta_{k}(v_{\boldsymbol{s}})\,p_{\boldsymbol{s}}(\theta,\lambda).$
@@
$\begin{split}A_{1}(\xi)\ &\leq\ \sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}\|\delta_{k}(v_{\boldsymbol{s}})\|_{X^{1}}p_{\boldsymbol{s}}(\theta,\lambda)\\ \ &\leq\ C\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}\|v_{\boldsymbol{s}}\|_{X^{1}}\sum_{k=0}^{k(\xi)}\,p_{\boldsymbol{s}}(\theta,\lambda)\\ \ &\leq\ C\log\xi\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}\|v_{\boldsymbol{s}}\|_{X^{1}}\,p_{\boldsymbol{s}}(\theta,\lambda).\end{split}$ (3.36)
@@
$\begin{split}\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}\|v_{\boldsymbol{s}}\|_{X^{1}}\,p_{\boldsymbol{s}}(\theta,\lambda)\ &\leq\ \left(\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}(\sigma_{1;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{1}})^{2}\right)^{1/2}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}p_{\boldsymbol{s}}^{2}(\theta,\lambda)\sigma_{1;{\boldsymbol{s}}}^{-2}\right)^{1/2}\\[6.45831pt] \ &\leq\ C\,\left(\sum_{\sigma_{1;{\boldsymbol{s}}}^{q_{1}}>\xi}p_{\boldsymbol{s}}^{2}(\theta,\lambda)\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\sigma_{1;{\boldsymbol{s}}}^{-(2-q_{1})}\right)^{1/2}\\[6.45831pt] \ &\leq\ C(\xi)^{-(1/q_{1}-1/2)}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}p_{\boldsymbol{s}}(2\theta,\lambda)\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\right)^{1/2}\\[6.45831pt] \ &\leq\ C\xi^{-(1/q_{1}-1/2)}.\end{split}$
This and (3.36) yield that
$A_{1}(\xi)\ \leq\ C\xi^{-(1/q_{1}-1/2)}\log\xi.$ (3.37)
We now give a bound for $A_{2}(\xi)$. Observe that $\vartheta\alpha/\tau=1/q_{1}-1/2$ and $\alpha/\tau=1-q_{2}/2$. Hence, employing (2.4), the assumption (3.18) and the Hölder inequality we get
$\begin{split}A_{2}(\xi)\ &\leq\ \sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\ \sum_{2^{k}>\left(\xi^{\vartheta}\sigma_{2;{\boldsymbol{s}}}^{-1}\right)^{1/\tau}}\|\delta_{k}(v_{\boldsymbol{s}})\|_{X^{1}}\,\big\|I_{\Lambda_{k}(\xi)\cap R_{\boldsymbol{s}}}\,H_{\boldsymbol{s}}\big\|_{L_{2}({\mathbb{R}}^{\infty},\gamma)}\\[6.45831pt] \ &\leq\ C\,\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\ \sum_{2^{k}>\left(\xi^{\vartheta}\sigma_{2;{\boldsymbol{s}}}^{-1}\right)^{1/\tau}}2^{-\alpha k}\|v_{\boldsymbol{s}}\|_{X^{2}}\,p_{\boldsymbol{s}}(\theta,\lambda)\\[6.45831pt] \ &\leq\ C\,\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\ \big(\sigma_{2;{\boldsymbol{s}}}^{-1/\tau}\xi^{\vartheta/\tau}\big)^{-\alpha}\|v_{\boldsymbol{s}}\|_{X^{2}}\,p_{\boldsymbol{s}}(\theta,\lambda)\\[6.45831pt] \ &\leq\ C\,\xi^{-\vartheta\alpha/\tau}\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\ \sigma_{2;{\boldsymbol{s}}}^{\alpha/\tau}\|v_{\boldsymbol{s}}\|_{X^{2}}\,p_{\boldsymbol{s}}(\theta,\lambda)\\[6.45831pt] \ &\leq\ C\,\xi^{-(1/q_{1}-1/2)}\ \sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\ \sigma_{2;{\boldsymbol{s}}}^{1-q_{2}/2}\|v_{\boldsymbol{s}}\|_{X^{2}}\,p_{\boldsymbol{s}}(\theta,\lambda)\\ \ &\leq\ C\,\xi^{-(1/q_{1}-1/2)}\ \left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}(\sigma_{2;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X^{2}})^{2}\right)^{1/2}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}p_{\boldsymbol{s}}^{2}(\theta,\lambda)\sigma_{2;{\boldsymbol{s}}}^{-2}\right)^{1/2}\\ \ &\leq\ C\,\xi^{-(1/q_{1}-1/2)}.\end{split}$
@@
$A_{2}(\xi)\ \leq\ C\xi^{-(1/q_{1}-1/2)}.$ (3.38)
Combining (3.37), (3.38) and (3.35) leads to the estimate
@@
We estimate the dimension of the space ${\mathcal{V}}(G(\xi))$. Put $q:=\tau q_{2}$ and define $q^{\prime}$ by $1/q^{\prime}+1/q=1$. Since $\alpha>1/q_{2}-1/2$, we have that $q>1$, consequently, by using the Hölder inequality and (3.18), we derive that
@@
$\xi_{n}^{-(1/q_{1}-1/2)}\asymp n^{-(1/q_{1}-1/2)\frac{\alpha}{\alpha+\delta}}\ =\ n^{-\beta}.$
@@
where $k_{n}:=\lfloor\log_{2}\xi_{n}\rfloor$ if $\alpha\leq 1/q_{2}-1/2$, and $k_{n}:=\lfloor\vartheta\tau^{-1}\log_{2}\xi_{n}\rfloor$ if $\alpha>1/q_{2}-1/2$, and for $k\in{\mathbb{N}}_{0}$ and $\xi>1$,
@@
Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that for the operator ${\mathcal{I}}^{*}_{G_{\rm ev}(\xi_{n})}:{\mathcal{L}}_{2}^{\mathcal{E}}(X^{2})\to{\mathcal{V}}(G_{\rm ev}(\xi_{n}))$, we have that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
@@
Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda(\xi_{n}))|\leq n$ and
@@
Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
@@ 4 Integration
    Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
@@
    Let $\phi\in(X^{1})^{\prime}$ be a bounded linear functional on $X^{1}$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $\dim{\mathcal{V}}(G_{\rm ev}(\xi_{n}))\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\|\phi\|_{(X^{1})^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (4.5)
@@
$\begin{split}\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v_{\rm ev}({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-{\mathcal{Q}}_{G_{\rm ev}(\xi_{n})}\langle\phi,v\rangle\right|\ &\leq\ \int_{{\mathbb{R}}^{\infty}}\left|\langle\phi,v_{\rm ev}({\boldsymbol{y}})-{\mathcal{I}}^{*}_{G_{\rm ev}(\xi_{n})}v_{\rm ev}({\boldsymbol{y}})\rangle\right|\,{\rm d}\gamma({\boldsymbol{y}})\\[6.45831pt] \ &\leq\ \int_{{\mathbb{R}}^{\infty}}\|\phi\|_{(X^{1})^{\prime}}\|v_{\rm ev}({\boldsymbol{y}})-{\mathcal{I}}^{*}_{G_{\rm ev}(\xi_{n})}v_{\rm ev}({\boldsymbol{y}})\|_{X^{1}}\,{\rm d}\gamma({\boldsymbol{y}})\\[6.45831pt] \ &\leq\ C\|\phi\|_{(X^{1})^{\prime}}\|v_{\rm ev}-{\mathcal{I}}^{*}_{G_{\rm ev}(\xi_{n})}v_{\rm ev}\|_{{\mathcal{L}}_{1}(X^{1})}\\[6.45831pt] \ &\leq\ C\|\phi\|_{(X^{1})^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}\end{split}$
@@
    Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
@@
    Let $\phi\in X^{\prime}$ be a bounded linear functional on $X$. Then there exists a constant $C$ such that for each $n\in{\mathbb{N}}$, there exists a number $\xi_{n}$ such that $|\Gamma(\Lambda_{\rm ev}(\xi_{n}))|\leq n$ and
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-Q_{\Lambda_{\rm ev}(\xi_{n})}\langle\phi,v\rangle\right|\ \leq\ C\|\phi\|_{X^{\prime}}n^{-(1/q-1/2)}.$ (4.8)
@@
Define for $\xi>1$,
$G_{2}(\xi):=\ \begin{cases}\big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{2}:\,2^{k}\sigma_{2;{\boldsymbol{s}}}^{q_{2}}\leq\xi\big\}\quad&{\rm if}\ \alpha\leq 1/q_{2}-1/2;\\ \big\{(k,{\boldsymbol{s}})\in{\mathbb{N}}_{0}\times{\mathbb{F}}_{2}:\,\sigma_{1;{\boldsymbol{s}}}^{q_{1}}\leq\xi,\ 2^{\tau k}\sigma_{2;{\boldsymbol{s}}}\leq\xi^{\vartheta}\big\}\quad&{\rm if}\ \alpha>1/q_{2}-1/2,\end{cases}$
where $\tau$ and $\vartheta$ are as in (3.6).
@@
Under the hypothesis of Theorem 3.1, assume additionally that the sequences $Y_{m}$, $m\in{\mathbb{N}}_{0}$, are symmetric.
@@
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{{\mathcal{Q}}}_{G_{2}(\xi_{n})}\langle\phi,v\rangle\right|\leq C\|\phi\|_{(X^{1})^{\prime}}\times\begin{cases}n^{-\alpha}&\text{if }\alpha\leq 1/q_{2}-1/2,\\ n^{-\beta}\log n&\text{if }\alpha>1/q_{2}-1/2.\end{cases}$ (4.10)
@@
    $\left|\int_{{\mathbb{R}}^{\infty}}\langle\phi,v({\boldsymbol{y}})\rangle\,\,{\rm d}\gamma({\boldsymbol{y}})-\tilde{Q}_{\Lambda_{2}(\xi_{n})}\langle\phi,v\rangle\right|\ \leq\ C\|\phi\|_{X^{\prime}}n^{-(1/q-1/2)}.$ (4.13)